Write an equation of the line that is parallel to the given line
y = x + 3 and the given point (5,0).
-----------------------------------------------------------------------

Write an equation of the line that is parallel to the given line
y = 2x + 3 and the given point (-4,1).

Answer by SwiftAlbatross(13)

y = x + 3 has a slope of 1 since the coefficient in front of x is 1. Therefore, the parallel line also has a slope of 1. Knowing this, we just need to find the y-intercept to put the parallel line in the form y = mx + b. Since we know the slope and we have the x and y coordinates of the given point which the line passes through, let us plug these into the y = mx + b equation to find the unknown y-intercept, b. So,
y = mx + b
0 = 1*5 + b
0 = 5 + b
-5 = -5
-5 = b

So the equation of the parallel line is y = x - 5. The other problem is very similar to this. I will just do the math.
y = mx + b
1 = 2*-4 + b
1 = -8 + b
+8 = +8
9 = b
y = 2x + 9